# Privacy Preservation Techniques (PPTs) in IoT Systems: A Scoping Review and Future Directions Author’s draft for soliciting feedback - February 20, 2025

### v1

@@ III Comparison with existing surveys and reviews
Paper PPT Review PPT Goals IoT applications Privacy Types PPT implementation in IoT computing layers Literature Review
[18] $\star$ $\star$ Systematic Mapping
[19] $\star$ $\star$ $\star$ Narrative Review
[20] $\star$ Systematic Mapping
[21] $\star$ Systematic Review
@@
: The study reviews the topic comprehensively,      $\star$: The study covers the topic partially,
@@
The comparative insights shown in Table I highlight the uniqueness of this scoping review. This study is the first to conduct a comprehensive scoping review of PPT in an IoT system, providing a detailed understanding of the technologies used in building PETs, IoT applications, privacy goals, the implementation of PPT across different layers of IoT computing, and the types of privacy associated with these techniques. We present an explanation of similar reviews related to this study as follows:
Kinza et al. [18] studied how privacy can be protected in IoT applications. They placed a particular emphasis on how fog computing can strengthen privacy in the IoT field. They discussed the PETs used for implementing privacy solutions: anonymization and data access control. Their study primarily focused on technical approaches such as homomorphic encryption, logistical scrambling, and distributive role-based access controls [18]. Additionally, the study clearly explained the types of privacy protection covered by the techniques they examined. While they reviewed many IoT applications, the specific privacy goals of the studied techniques received relatively limited attention.
Abi Sen et al. [19] partially discuss some technologies of use in their study, primarily focusing on mathematical models and cryptography used for privacy preservation in IoT systems in the literature. However, their study covers common privacy goals and IoT applications (except newly introduced IoT applications in multimedia, social networks, and wearables). They primarily emphasized the preservation of user and location privacy as the main focus of their review, highlighting the challenges associated with these specific privacy types in their study [19]. Additionally, they explain the significance of central cloud storage in IoT and its inherent limitations, proposing the adoption of Fog computing to address these limitations through decentralized storage and processing capabilities [19].
Torre et al. [20] addressed a few specific privacy goals of PPTs in their study. These goals include anonymization, access control, data flow privacy, data minimization, and obfuscation. While the authors [20] provide some insights into the technology used for building PETs, their primary focus centers on cryptography in this category. However, they also touch upon blockchain and machine learning as pertinent approaches. Furthermore, their review predominantly focuses on IoT applications in health, transport, and wearable [20] without covering the recently introduced IoT applications.
@@
Some gaps were observed in previous reviews on PPTs in the IoT environment, especially in the technology used to build PETs, privacy goals, privacy preservation types, the IoT application of interest, and the PPT implementation in the IoT computing layer. Prior reviews on PPT in IoT focus on studies that use technical approaches to achieve PPT. However, PPTs in IoT should also incorporate non-technical approaches such as informed consent.
Furthermore, this study discovered that no prior reviews have covered new trends in IoT applications in areas such as social networks and multimedia.
@@
This study addressed some of these research gaps by discussing potential solutions (see Section X) based on specific research questions. This study also explored the goals of various privacy techniques, including hybrid solutions that combine multiple approaches. Additionally, this study focused on different types of privacy, especially those that have received less attention from researchers, such as device identity privacy, and analyzed IoT application areas, particularly those recently introduced in IoT.
IV The Five-step Research Methodology
@@
This study adopted PRISMA-ScR (Preferred Reporting Items for Systematic Reviews and Meta-Analyses Extension for Scoping Reviews) due to its comprehensive search strategy, which helps map the literature into meaningful categories within a broad context. The scoping review reporting guideline was initially introduced by Tricco et al. [8]. This guideline features a detailed checklist and explanations of PRISMA-ScR. This study implements PRISMA-ScR within the framework established by Arksey et al. [39]. Arksey’s framework consists of five stages: (1) identifying research questions, (2) identifying relevant studies, (3) selecting pertinent studies, (4) charting the collected data (data visualization), and (5) synthesizing and presenting the findings. It is important to note that this guideline inherently includes the essential components outlined in the PRISMA reporting methodology.
V Step 1: Identifying the research questions
@@
To better understand the factors affecting privacy and its preservation in IoT systems, we formulated essential questions not addressed in previous studies.
@@
These inquiries uncover what, why, and how certain IoT activities and applications influence the techniques employed to preserve privacy in IoT systems. Here, the study presents a set of research questions (RQs) that are not adequately addressed in prior reviews on PPTs in the IoT.
  • RQ1: What technologies are prominently used to develop PETs in IoT systems?
  • RQ2: What are the dominant PPT’s privacy goals in IoT applications?
  • RQ3: What are the commonly known and recent IoT applications added to the literature?
  • RQ4: What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT most implemented from the literature?
  • RQ5: What privacy types are the main focus of most PPTs in IoT systems?
VI Step 2: Identifying relevant studies
@@
To understand privacy preservation in the IoT environment, this study queried selected databases with the following search expression: [(”Privacy preservation techniques” OR ”Privacy protection techniques” OR ”Privacy preservation methods” OR ”Privacy protection Approaches” OR ”Privacy protection methods” OR ”Privacy protection techniques”) AND (”Internet of Things” OR ”IoT” OR ”Smart Things” OR ”Smart environment” OR ”Internet of Everything”)].
@@
This study relied on reputable sources from seven databases: IEEE Explorer, ACM Digital Library, ScienceDirect, SpringerLink, Sage Journal, IET, and Taylor and Francis, ensuring the robustness and reliability of our findings. As illustrated in Figure 2, a total of 725 papers were initially collected, including conference proceedings, journal articles, white papers, websites, and theses. The first step of this study involved reviewing the titles, during which 223 papers were excluded for reasons such as being white papers, surveys, non-academic documents, or irrelevant content. Next, the abstracts and conclusions of the remaining 502 papers were examined, excluding 245 papers that were not related to the Internet of Things (IoT). From the remaining 257 papers, twelve (12) papers that lacked privacy considerations were excluded, and two duplicates were removed. This process resulted in a final total of 243 papers for our review.
@@
This study applied both inclusion and exclusion criteria. The inclusion criteria required papers to focus on PPT in IoT, propose at least one PPT for IoT, be peer-reviewed, and be published between 2010 and 2023. However, the exclusion criteria ruled out secondary manuscripts, non-IoT-related papers, non-educational papers, white papers, non-scientific documents, and non-English papers.
VII Step 3: Selection of PPT publications in IoT
@@
This study explored seven databases to examine their content on PPTs in IoT systems. IEEE Explorer was the most prolific, contributing about 59% of the total publications. In contrast, Taylor and Francis, Sage and IET had the lowest contributions, as shown in Figures LABEL:fig:Journal-Database-bar and LABEL:fig:Journal-Database-pie.
The volume of PPT publications declined from 2013 to 2017, reached its peak in 2019, and has been increasing again from 2019 to 2023, with another notable peak occurring in 2021. By the first quarter of 2023, which marked the end of our paper collection, 42 published papers had been gathered that year, indicating a significant rise in PPT publications, as illustrated in Figure 4.
@@
Figure 5 shows that IEEE Explorer houses the highest number of PPT papers based on PPT goals in IoT systems, followed by ScienceDirect. Taylor and Francis, IET, and Sage contributed modestly, each offering between one and two PPT papers in IoT.
@@
Figure 4: Number of publications on PPTs in IoT per year between 2010-2023
Figure 5: Papers on PPT goals in IoT system and their published Database
VIII Step 4: Charting the data collected on PPT publications
The collected papers are categorized by PPT goals, technology used for building PET, IoT applications utilizing PPTs, computing layers where PPT is integrated within the IoT architecture, and type of privacy protection. Our results will address our research questions in subsection IX by reviewing each category.
VIII-A PPT and PPT Goals
@@
VIII-A1 Anonymity
Anonymity in data privacy involves altering the original data format to hide sensitive parts of the that could lead to the identification of the data subject [40, 41, 42]. Various techniques are used to achieve anonymity:
@@ VIII-A1 Anonymity
  • Data Generalization: Reducing data precision to a level that prevents easy linkage to individuals. It is used in maintaining data accuracy for tasks like medical data analysis [44, 45].
@@
  • Data Swapping: Exchanging data attributes value with values from different records to preserve statistical characteristics while obscuring individual data records[48] [49].
@@
VIII-A2 Data flow privacy
@@
  • Data Sharing protection: Data sharing is essential in IoT systems but poses privacy risks. Different PPTs are employed to protect sensitive data during transmission and sharing processes [52].
@@
  • Data verification protection: This process reviews and helps improve existing privacy implementation in an IoT system. Techniques such as Petri net analysis proposed by Malek et al. [54] are used to analyze and improve privacy implementations in IoT systems, identifying and addressing privacy violations.
  • Route path protection: This is crucial in location-based services in IoT networks. For example, obfuscation of path is proposed by Yang et al. [55] to protect users’ path privacy.
  • Collaborative Privacy protection: Collaboration in distributed architecture is essential. However, trust needs to be established among the entities involved. Therefore, privacy techniques that collectively protect data sharing during collaboration are important to establish trust among the collaborative entities in the IoT environment [56].
VIII-A3 Data Access Control
@@
  • Blockchain-Based Access Control (BBAC): Access control based on blockchain technology includes transaction-based and smart contract-based mechanisms. Transaction-based control allows granting, revoking, or delegating access, while smart contracts enforce access policies [57].
@@ VIII-A5 Data Isolation
It is the process of confining IoT devices or services away from other network segments. Thapa et al. [61] describe different levels of isolation, and Zhang et al. [62] proposed virtualization and containerization in smart healthcare systems.
VIII-A6 Data Minimization
@@
Reduces data collection and processing in IoT systems to ensure only relevant data are used, preventing over-utilization. Studies focus on data minimization in IoT include [63, 64, 65].
VIII-A7 Multiple Goals
@@
Some PPTs aim to achieve multiple goals simultaneously: Data anonymization and access control in [66, 67, 68], data flow privacy and access control in [69], data flow privacy and minimization in [70], data isolation and distortion in [27]. Figure II presents other papers that achieved multiple goals in their proposed work.
@@
Our observations: From our study, anonymity emerges as the primary goal of PPTs across IoT application systems, with data masking being the most commonly used method to achieve this. This observation is supported by our findings as shown in Table III. Additionally, many PPTs address multiple privacy goals simultaneously, as shown in Table II. These findings underscore the importance of achieving anonymity in protecting sensitive data across diverse IoT applications while highlighting the need for comprehensive privacy strategies that simultaneously address multiple facets of data protection.
@@
Anonymity and Access control [66] [67] [68] [71] [72]
Data flow protection and Data minimization [70]
Data minimization and Data Isolation [73]
Anonymity and Data isolation [27] [74] [75] [76] [77] [78]
Data flow protection and Access control [69]
@@
Anonymity and Data flow protection [80] [81]
Access control and Data Minimization [82] [83]
Table II: Papers on PPT with multiple privacy goals
{tblr}
width = colspec = Q[120]Q[120]Q[320]Q[70], cell21 = r=5, cell71 = r=5, cell121 = r=7, vlines, hline1-2,7,12,20-23 = -, hline3-6,8-11,13-19 = 2-4, PPT Goals & PPT method Paper Number of Paper
Anonymity Data masking [84] [85] [86] [87] [88] [89] [90] [91] [92] [93] [94] [95] [96] [97] [98] [99] [100] [101] [102] [103] [104] [105] [106] [107] [108] [109] [110] [111] [112] [113] [114] [115] [116] [117] [118] [119] [120] [121] [122] [123] [124] [125] [126] [127] [128] [129] [130] [131] [132]   49
Data generalization [133] [134] [135] [136] [137] [138] [139] [140] [141] [142] [143] [144] [145] [146] [147] [148] [149] 17
Data perturbation  [32] [150], [151] [152] [153] [154] [155] [156] [157] [158] [159] [160] [161] [162] [163][164] [165] [166] [167] [168] [25] [169] [170] [171] [172] [24] [173] [174] [175] [176] [177] [178] [179] [180] [181] [182] [183] [184] 38
Data Pseudonymization [185] [186] [187] [188] [189] [190] [191] [192] [193] [194] [195] [23] [196] [197] [198] [199] [200] [201] [202] [203] 20
Data Swapping [204] 1
Data Flow Privacy Data Sharing protection [205] [28] [206] [207] [208] [209] [210] [211] [212] [213] [214] [215] [216] 13
Multi-path routing [217] [218] [219] 3
Data Privacy Verification [220] [221] 2
Route path protection [222] [223] [224] [225] [226] [227] [228] [229] [230] [231] 10
Collaborative PP [232] [233] [234] [235] 4
Access Control RBAC [236] [237] [238] [239] 4
ABAC [240] [241] [242] [26] [243] [244] [245] [246] [247] [248] [249] 11
RBAC and ABAC [250] 1
BBAC [251] [252] [253] [22] [254] [255][256] [257] [258] [259] [260] [261] [262] [263] [264] [265] [266] [267] 18
CapBAC [268] [269] [270] 3
UCON [271] [272] [273] 3
ReBAC [274] 1
DAC [275] [276] [277] 3
Data Privacy Awareness - [278] [279] [280] [281] [282] 5
Data Isolation - [283] [284] [285] [286] [287] 4
Data Minimization - [288] [289] [290] [291] [292] [293] [294] [295] [296] [297] [298] [299] 12
Table III: Papers on PPT Goals and their Approaches
VIII-B Technology used in building PETs

### v2

@@ 3 Comparison of this paper with other surveys and review papers
Paper PPT Review PPT Goal IoT application Privacy Type PPT Implementation in IoT Computing Layers Literature Review
[431] Systematic Mapping
[6] Narrative Review
[509] Systematic Mapping
[15] Systematic Review
@@
: The study reviews the topic comprehensively,      : The study covers the topic partially,
@@
Kinza et al. [431] conducted a study on how privacy can be protected in IoT applications. They placed a particular emphasis on how fog computing can strengthen privacy in the IoT field. They discussed the technology used for implementing privacy solutions which are anonymization and data access control. Their study primarily focused on technical approaches, such as homomorphic encryption, logistical scrambling, and distributive role-based access controls [431]. Additionally, the researchers clearly explained the types of privacy protection covered by the techniques they examined. While they reviewed many IoT applications, the specific privacy objectives of the studied techniques received relatively limited attention.
In their review, Abi Sen et al. [6] partially discuss some technologies of use in their study, primarily focusing on mathematical models and cryptography used for privacy preservation in IoT systems in the literature. However, their study covers common privacy goals and IoT applications (except newly introduced IoT applications in multimedia, social networks, and wearables). They primarily emphasized the preservation of user and location privacy as the central focus of their review, highlighting the challenges associated with these specific privacy protection types in their study.[6]. Additionally, they explain the significance of central cloud storage in IoT and its inherent limitations, proposing the adoption of Fog computing to address these limitations through decentralized storage and processing capabilities.
Torre et al. [509] address a few specific privacy goals of privacy techniques reviewed in their study. These goals include anonymization, access control, data flow privacy, data minimization, and obfuscation. While the authors [509] provide some insights into the technology used for privacy preservation, their primary focus centers on cryptography. However, they also touch upon blockchain and machine learning as pertinent approaches. Furthermore, their review predominantly focuses on IoT applications in health, transport, and wearable [509].
@@
Some gaps were observed in previous reviews on PPTs in the IoT environment, especially in the technology used to build PETs, privacy goals, privacy preservation types, the IoT application of interest, and the PPT implementation in the IoT computing layer. Prior reviews on PPT in IoT focus on studies that use technical approaches to achieve PPT. However, PPTs in IoT should also incorporate non-technical approaches such as informed consent. Furthermore, this study discovered that no prior reviews have covered new trends in IoT applications in areas such as social networks and multimedia.
@@
We addressed some of the research gaps by discussing potential solutions (see section 10) in response to specific research questions in section 5. Our study also examined the objectives of various privacy techniques, including hybrid solutions that integrate multiple approaches. Additionally, we focused on various types of privacy, particularly those that have received less attention from researchers, such as device identity privacy, and analyzed IoT application areas, specifically those recently introduced in the context of IoT.
4 The Five-step Research Methodology
@@
We adopted PRISMA-ScR (Preferred Reporting Items for Systematic Reviews and Meta-Analyses Extension for Scoping Reviews) due to its comprehensive search strategy, which helps map the literature into meaningful categories within a broad context. The scoping review reporting guideline was initially introduced by Tricco et al. [510]. This guideline features a detailed checklist and explanations of PRISMA-ScR. This study implements PRISMA-ScR within the framework established by Arksey et al. [57]. Arksey’s framework consists of five stages: (1) identifying research questions, (2) identifying relevant studies, (3) selecting pertinent studies, (4) charting the collected data (data visualization), and (5) synthesizing and presenting the findings. It is essential to note that this guideline inherently encompasses the key components outlined in the PRISMA reporting methodology.
5 Step 1: Identifying the research questions
@@
To better understand the factors that affect privacy and its preservation in IoT systems, we formulated essential questions that have not been addressed in previous studies.
@@
These inquiries reveal what, why, and how certain IoT activities and applications impact the techniques used to preserve privacy in IoT systems. Here, the study presents a set of research questions (RQs) that are not adequately addressed in prior reviews on PPTs in the IoT. These research questions are (1) What technologies are prominently used to develop PETs in IoT systems? (2) What are the dominant PPT’s privacy goals in IoT applications? , (3) What are the commonly known and recent IoT applications added to the literature?, (4) What computing layer in IoT architectures are PPTs mostly implemented within IoT systems, and which IoT application has PPT most implemented from the literature?,(5) What privacy types are the main focus of most PPTs in IoT systems?
6 Step 2: Identifying relevant studies
@@
To understand privacy preservation in the IoT environment, we queried selected databases with the following search expression: [(”Privacy preservation techniques” OR ”Privacy protection techniques” OR ”Privacy preservation methods” OR ”Privacy protection Approaches” OR ”Privacy protection methods” OR ”Privacy protection techniques”) AND (”Internet of Things” OR ”IoT” OR ”Smart Things” OR ”Smart environment” OR ”Internet of Everything”)].
@@
Our work relied on reputable sources from seven databases: IEEE Xplore, ACM Digital Library, ScienceDirect, SpringerLink, Sage Journal, Wiley Online, and Taylor & Francis, ensuring the robustness and reliability of our findings. As illustrated in Figure 5, a total of 1177 papers were initially collected, including conference proceedings, journal articles, white papers, websites, and theses. The first step of this study involved reviewing the titles, during which 502 papers were excluded for reasons such as being white papers, surveys, non-academic documents, or articles with irrelevant content. Next, the abstracts and conclusions of the remaining 675 papers were examined, excluding 327 papers that were not related to the Internet of Things (IoT). From the remaining 348 papers, twelve (15) papers that lacked privacy considerations were excluded, and four duplicates were removed. This process resulted in a final total of 329 papers for our review.
@@
We applied both inclusion and exclusion criteria. The inclusion criteria required papers to focus on PPT in IoT, propose at least one PPT for IoT, be peer-reviewed, and be published between 2010 and 2025. However, the exclusion criteria ruled out secondary manuscripts, non-IoT-related papers, non-educational papers, white papers, non-scientific documents, and non-English papers.
7 Step 3: Selection of PPT publications in IoT
@@
Seven databases were investigated to gather publications relevant to PPTs in IoT systems. IEEE Xplore was the most prolific, contributing about 46% of the total publications. In contrast, Taylor & Francis, Sage, and Wiley Online had the lowest contributions, as shown in Figures 6 and 7.
The volume of PPT publications collected increased from 2010 to 2018, reached its peak in 2019, but declined in 2020 and 2022, with another notable peak in 2021. In 2023 and 2024, there were extensive collections of papers; however, early December of 2025 marked the end of our paper collection, which had fewer publications than in 2024. We expect more publications to be released before the end of 2025, as illustrated in Figure 8.
@@
Figure 9 shows that IEEE Xplore houses the highest number of PPT papers based on PPT goals in IoT systems, followed by ScienceDirect. Taylor & Francis, Wiley Online, and Sage contributed modestly, each offering between one and two PPT papers in the field of IoT.
@@
7.1 Selection of PPTs publication in IoT
Our study examined seven(7) databases to explore their content on PPTs in the IoT system. Among these databases, IEEE Explorer emerged as the most prolific, boasting approximately 59% of the total publications on the subject. However, Taylor and Francis, Sage, and IET exhibited the lowest contribution, each accounting for a mere 0% of the publications encompassed in our review, as illustrated in both Figure 6 and 7.
To adhere to our exclusion criteria, we confined our focus to journal reviews published between 2010 and 2023. Notably, the early years from 2013 to 2017 witnessed a dip in publications, with 2019 marking the zenith in publication volume. Subsequently, from 2019 to 2023, it experienced a publication surge, with another notable peak in 2021. While collecting journals in April 2023, we amassed 31 journals published in 2023, indicating a noteworthy increase in journal publications on PPTs in the first quarter of that year, as highlighted in Figure 8.
Furthermore, Figure 9 elucidates that the IEEE Explorer database houses the highest number of journals on PPTs categorized based on PPT goals in IoT systems, with ScienceDirect following closely as the second-highest database featuring an extensive collection of journals on PPTs in IoT. In contrast, Taylor and Francis, The Institute of Engineering and Technology (IET), and Sage contributed modestly, each offering between one and two journals on PPTs in IoT.
Figure 6: Number of papers in Database on PPTs in IoT Applications considered in our review
Figure 7: Percentage of papers in Database on PPTs in IoT Applications considered in our review
Figure 8: Number of publications on PPTs in IoT per year between 2010-2025
Figure 9: Papers on PPT goals in IoT systems and their published Database
8 Step 4: Charting the data collected on PPT publications
The collected papers are categorized by PPT goals, technology used for building PET, IoT applications utilizing PPTs, computing layers where PPT is integrated within the IoT architecture, and the type of privacy protection. Our results address our research questions in subsection 9 by reviewing each category as follows.
8.1 PPT and PPT Goals
@@
8.1.1 Anonymity
Anonymity in data privacy involves altering the original data format to hide sensitive parts of the data that could lead to the identification of the data subject [166, 338, 421]. Various techniques are used to achieve anonymity:
@@ 8.1.1 Anonymity
  • Data generalization: Reducing data precision to a level that prevents easy linkage to individuals. It is used to maintain data accuracy for tasks such as medical data analysis [533, 115].
@@
  • Data swapping: Exchanging data attributes’ values with values from different records to preserve statistical characteristics while obscuring individual data records[191] [330].
@@
8.1.2 Data Flow Privacy
@@
  • Data sharing protection: Data sharing is essential in IoT systems, but it also poses significant privacy risks. Different PPTs are employed to protect sensitive data during transmission and sharing processes [74].
@@
  • Data verification protection: This process reviews and helps improve existing privacy implementation in an IoT system. Techniques such as Petri net analysis, proposed by Malek et al. [1], are used to analyze and improve privacy implementations in IoT systems, identifying and addressing privacy violations.
  • Route path protection: This is crucial in location-based services in IoT networks. For example, Yang et al. [581] propose obfuscating the path to protect users’ path privacy.
  • Collaborative privacy protection: Collaboration in a distributed architecture is essential. However, trust needs to be established among the entities involved. Therefore, privacy techniques that collectively protect data sharing during collaboration are crucial for establishing trust among collaborative entities in the IoT environment [610].
8.1.3 Data Access Control
@@
  • Blockchain-Based Access Control (BBAC): Access control based on blockchain technology includes transaction-based and smart contract-based mechanisms. Transaction-based control enables the granting, revoking, or delegating of access, while smart contracts enforce access policies [395].
@@ 8.1.5 Data Isolation
This process involves confining IoT devices or services within their own network segment. Thapa et al. [504] describe different levels of isolation, and Zhang et al. [621] proposed virtualization and containerization in smart healthcare systems.
8.1.6 Data Minimization
@@
Reduces data collection and processing in IoT systems to ensure only relevant data is used, preventing over-utilization. Studies focus on data minimization in IoT include [4, 467, 363].
8.1.7 Multiple Goals
@@
Some PPTs aim to achieve multiple goals simultaneously: Data anonymization and access control in [384, 587, 10], data flow privacy and access control in [373], data flow privacy and minimization in [122], data isolation and distortion in [422]. Figure 3 presents other papers that achieved multiple goals in their proposed work.
@@
Our observations: From our study, anonymity emerges as the primary goal of PPTs across IoT application systems. Approximately 52% of publications employ anonymity, with data perturbation the main privacy goal, followed by data masking and data pseudonymization. This observation is supported by our findings, as shown in Table 2. Additionally, many PPTs address multiple privacy goals simultaneously, as shown in Table 3. We observed an increase in the use of the PPT approach targeting multiple goals. The number of publications that achieve both anonymity and data flow protection increased from 2023 to 2025, as shown in Table 3. These findings underscore the importance of achieving anonymity in protecting sensitive data across diverse IoT applications while highlighting the need for comprehensive privacy strategies that simultaneously address multiple facets of data protection.
Table 2: Papers on PPT Goals and their Approaches
PPT Goals PPT approach Paper Number of Paper
Anonymity Data masking [183, 563, 173, 185, 346, 553, 93, 435, 125, 543, 144, 583, 530, 623, 641, 100, 412, 329, 13, 462, 520, 598, 47, 294, 270, 90, 550, 496, 619, 557, 612, 371, 532, 535, 150, 595, 324, 486, 283, 145, 17, 578, 580, 61, 378, 566, 390, 518, 586, 170, 89, 236, 228, 257, 406, 134, 564, 593, 320, 200, 76, 123, 452, 314, 528, 627, 215, 360, 261, 468, 317, 318, 466] 73
Data generalization [287, 506, 310, 427, 465, 364, 503, 472, 167, 214, 570, 639, 358, 491, 210, 382, 233, 148, 428, 277, 309, 240, 42] 23
Data perturbation [170, 51, 628, 30, 304, 579, 562, 203, 508, 470, 218, 591, 80, 531, 387, 313, 97, 622, 556, 433, 194, 551, 187, 196, 561, 87, 285, 297, 576, 29, 614, 31, 560, 153, 546, 5, 611, 149, 325, 463, 263, 458, 9, 352, 64, 18, 481, 296, 574, 243, 3, 517, 464, 50, 102, 388, 248, 179, 238, 251, 281, 222, 620, 479, 128, 168, 198, 339, 523, 230, 75, 519, 590, 624, 385, 419, 337, 495, 497, 96, 81, 630, 609, 282, 262] 85
Data Pseudonymization [356, 21, 489, 202, 402, 223, 119, 38, 32, 353, 615, 439, 84, 386, 542, 141, 484, 480, 23, 441, 194, 618, 265, 326, 547, 603, 293, 349, 177, 327, 540, 157, 258, 127, 425, 588, 126, 292, 552, 424, 20, 136, 537, 268, 369, 295, 342, 584, 477, 638, 189, 456, 48, 229, 592, 286, 527, 118, 181, 528, 526, 322, 609, 147, 380, 444, 117, 65, 137, 516, 68, 341] 72
Data Swapping [575] 1
Data Flow Privacy Data Sharing protection [432, 246, 120, 199, 453, 231, 398, 289, 549, 529, 332, 77, 522, 146, 411, 174, 158, 632, 407, 483, 321, 152, 274, 260, 83, 503, 333, 505, 451, 455, 103, 541, 367, 227, 536, 404, 335, 405, 211, 106, 347, 474, 114, 2, 636, 303, 316, 49, 429, 224, 443, 256, 271, 596, 312, 376, 609, 98, 379, 73, 116, 101, 104] 63
Multi-path routing [171, 205, 449] 3
Data Privacy Verification [589, 613, 14, 454, 607, 567, 291, 197, 394] 9
Route path protection [569, 460, 461, 577, 300, 392, 485, 267, 204, 206, 336, 43, 165, 476, 555] 15
Collaborative PP [253, 389, 511, 201, 383, 399, 546, 160, 357, 40, 255, 140, 186, 82, 490, 28, 585, 559, 539, 445, 617, 442, 528, 403, 276, 192, 55, 159] 28
Access Control RBAC [500, 340, 417, 446, 558, 78, 420, 244] 8
ABAC [604, 69, 348, 272, 599, 111, 629, 374, 637, 499, 471, 414, 469, 507, 355, 143, 161, 269, 568, 640] 20
RBAC and ABAC [221] 1
BBAC [434, 121, 72, 264, 154, 85, 306, 308, 305, 172, 644, 275, 302, 319, 397, 416, 7, 473, 193, 135, 514, 45, 207] 23
CapBAC [408, 232, 44, 164] 4
UCON [86, 169, 94, 133] 4
ReBAC [370] 1
DAC [298, 354, 217] 3
Data Privacy Awareness - [372, 634, 41, 438, 501, 410, 33, 34, 548, 372, 631, 108] 12
Data Isolation - [334, 247, 278, 393, 440, 138, 266, 91, 249, 448, 95, 107, 139, 646, 544, 113, 26, 16, 377] 18
Data Minimization - [565, 512, 626, 635, 46, 601, 436, 19, 423, 162, 79, 572, 112, 54, 409, 212, 37, 226, 343, 400] 20
Table 3: Papers on PPT with multiple privacy goals
@@
Anonymity and Data flow protection [418, 391, 597, 252, 273, 625, 250, 549, 498, 182, 27, 239, 616, 11, 290, 105, 118, 188, 618, 521, 130, 457, 67, 515, 88, 56, 513, 415, 594, 482, 401, 279, 571]
Access control and Data Minimization [237, 175]
Anonymity and Data Privacy Awareness [216]
Anonymity and Access control [384, 587, 10, 381, 92, 331, 217, 608, 554, 179]
Anonymity and Data isolation [422, 225, 235, 573, 643, 351, 129, 538]
Anonymity and Data Minimization [237, 175, 234, 488, 517, 315, 525, 259]
Data flow protection and Data minimization [122, 8, 605, 242, 53]
Data minimization and Data Isolation [241, 534, 450]
Data Flow Privacy and Access control [373, 524, 307, 58, 12, 582, 288]
Data Flow Privacy and Data Isolation [396, 220]
Anonymity, Access Control and Data Flow Privacy [344, 284]
Anonymity, Data Isolation and Data Flow Privacy [39]
8.2 Technology used in building PETs